# Implicit Dynamical Flow Fusion (IDFF) for Generative Modeling

### v1

@@ 3.2 IDFF time-series adaptation and sampling process
[…]
@@
    if $K==1$ then
         $k=1$
         ${\mathbf{x}}_{1}\sim p_{1}({\mathbf{x}})$, ${\mathbf{x}}_{0}\sim p_{0}({\mathbf{x}})$
    else
         $k\sim\{1,...,K\}$ with equal probability
         ${\mathbf{x}}_{1}\sim p_{1}({\mathbf{x}}^{k})$, ${\mathbf{x}}_{0}\sim p_{1}({\mathbf{x}}^{k-1}_{1})$
    end if
@@
    ${\mathbf{x}}\sim\mathcal{N}(\bm{\mu_{t}},\sigma_{t}^{2}\textbf{I})$
    $\mathcal{L}_{\text{IDFF}}(\theta)\leftarrow\beta(t)^{2}\|\hat{{\mathbf{x}}}_{1}({\mathbf{x}},t,k;\theta)-{\mathbf{x}}_{1}\|^{2}+\lambda(t)^{2}\left\|\bm{\epsilon}({\mathbf{x}},t;\theta)-\nabla_{x}\log p_{t}({\mathbf{x}}|{\mathbf{x}}_{1})\right\|^{2}$
@@
for $k$ in $\{1,...,K\}$ do
    if $k==1$ then
@@
    else
         ${\mathbf{x}}^{k}_{0}\sim\mathcal{N}({\mathbf{x}}^{k-1}_{1},\sigma_{0}\textbf{I})$
    end if
@@
         ${\mathbf{w}}^{k}_{t}\leftarrow\sqrt{(1-\sigma^{2}_{t})}\frac{\hat{{\mathbf{x}}}_{1}({\mathbf{x}},t,k;\theta)-{\mathbf{x}}^{k}_{t}}{1-t}+\frac{2\gamma-1}{2}\sigma_{t}\hat{\epsilon}({\mathbf{x}},t,k;\theta)$
@@
end for
return Samples $\{{\mathbf{x}}^{k}_{1}\}_{k=1}^{k=K}$
@@ 4 Related Work
Our proposed IDFF model extends state-of-the-art generative modeling techniques, particularly diffusion models such as denoising diffusion probabilistic models (DDPMs) (Ho et al., 2020) and CFMs. DDPMs rely on SDEs and Langevin dynamics for sampling (Yan et al., 2021), while CFMs use deterministic ODE solvers to sample from learned vector fields. Both methods are effective at generating high-quality samples but require many iterations, leading to computational inefficiency.
To address these limitations, denoising diffusion implicit models (DDIMs)(Song et al., 2020b) were introduced, offering a more efficient sampling process by deterministically generating samples from latent variables. Unlike DDPMs, which rely on Langevin dynamics, DDIMs employ a variational approach that accelerates sampling while preserving output quality.
In the realm of CFMs, recent advancements include Schrödinger-based methods (Richter et al., 2023) and conjugate/splitting-based integrators (Pandey et al., 2023), which aim to improve sampling efficiency. However, these methods still struggle to significantly reduce the NFEs, a key factor for computational efficiency. Other approaches to enhance sampling in diffusion-based generative models include the generalized Schrodinger bridge, which seeks stochastic evolution between prior and target distributions (Richter et al., 2023), and distillation frameworks like Flash (Kohler et al., 2024), which offer diverse sample generation in fewer steps by addressing training-inference discrepancies.
IDFF builds upon these developments and introduces several key improvements. Rather than directly learning vector fields like CFMs, IDFF focuses on learning the evolving samples $\bm{x}_{t}$ at each time step. It implicitly constructs the vector field during sampling through a reparameterization approach (Eq. 10). By incorporating a momentum term and a scoring model, IDFF allows for larger sampling steps, reducing NFE by up to a factor of 10 compared to CFMs, as illustrated in Figure 2.
A distinguishing feature of IDFF is its ability to jointly learn an implicit flow from the generated samples alongside the scoring model, setting it apart from models like DDIM, which focuses solely on learning a scoring function. By leveraging OT to couple samples between $t=0$ and $t=1$, IDFF accelerates convergence during training while maintaining low NFE and improving sample quality. It is evident from our image generation experiments (Tables 1 and 2). In comparisons with DDPM, DDIM, and DPM-Solver (another model that improves the NFEs for diffusion models)Lu et al. (2022a), IDFF consistently demonstrates superior efficiency and quality.
In conclusion, IDFF integrates the strengths of implicit flow and score-based learning, offering a novel and efficient framework for generative modeling that significantly reduces computational costs while producing high-quality samples.
5 Experiments
@@
We conducted two general classes of experiments to demonstrate the effectiveness of IDFF in generative modeling on static data, such as images and time-series data, including simulated chaotic systems (e.g., attractors), molecular dynamics, and sea surface temperature forecasting.
@@ 5.1 Image Generation
In this section, we showcase the superior performance of IDFF compared to CFMs in image generation, especially in the number of function evaluations (NFE). We trained IDFF models on datasets such as CIFAR-10 and CelebA (additional results for ImageNet-64, CelebA-HQ, LSUN Bedrooms, and LSUN Church are available in Appendix A.2). We utilized a CNN-based UNet (ADM) (Dhariwal and Nichol, 2021) to simultaneously model both $\hat{{\mathbf{x}}}_{1}(.,t;\theta)$ and $\bm{\epsilon}(.,t;\theta)$ from Equation 11 for all image data except ImageNet-64. For ImageNet-64, we employed a DiT architecture (’DiT-L/4’)(Peebles and Xie, 2022). To implement this, we doubled the input channels of the networks and fed the augmented input, $({\mathbf{x}}_{t},{\mathbf{x}}_{t})$, into the network. The outputs were then split into $(\hat{{\mathbf{x}}}_{1}(.,t;\theta),\bm{\epsilon}(.,t;\theta))$.
@@
To ensure a fair comparison with other models, we used commonly employed metrics to evaluate CFMs’ performance, including negative log-likelihood (NLL) measured in bits per dimension (BPD), sample quality as measured by the Frechet Inception Distance (FID), and the average NFE required for the reported FID and NLL, averaged over 50k samples, to assess the computational efficiency of our method. We generate 50k images to calculate the FID against the corresponding reference statistics of each dataset. All models use the same architecture, with specific details provided in Appendix A.1.4.
We also examined the effect of $\sigma_{0}$ on sample quality for both CIFAR-10 and CelebA datasets in Appendix A.1.3, finding that a value of $\sigma_{0}=0.6$ achieved the best FID for both datasets, as shown in Table 1 and Table 2. IDFF achieves comparable FID scores with ten times fewer NFEs than CFM and DDPM on the CIFAR-10 and CelebA datasets. IDFF’s qualitative results are provided in Figure 1 (additional samples are available in Appendix A.2).
@@ 5.2.1 Molecular Dynamics Simulation
Molecular dynamics simulation is a crucial tool in quantum mechanics for understanding the dynamics of molecular behavior at the atomic level (Cazorla and Boronat, 2017). Here, we simulated a fully extended polyalanine structure for 400 picoseconds in a vacuum environment at a temperature of $T=300K$.
@@
This molecular structure consists of 253 atoms with 46 dihedral angles. We trained an IDFF model to accurately generate the dihedral angles for the polyalanine structure from scratch. The distributions of the actual and generated dihedral angles for all 46 angles are presented in Figure 3 (A) and (B), respectively. Figure 3 (C) illustrates the dihedral angles for one of the alanine molecules and its associated dihedral angles. A complete trajectory showcasing a pair of actual and generated dihedral angles is depicted in Figure 3 (D).
To assess the performance of IDFF in the MD simulation, we utilized several evaluation metrics, including root mean squared error (RMSE), mean absolute error (MAE), and correlation coefficients (CC) between the generated and actual trajectories of dihedral angles. We benchmarked IDFF against well-known dynamical models for time series, including dynamical VAE models such as SRNNs (Fraccaro et al., 2016), VRNNs (Chung et al., 2015), and Neural ODEs (NODEs) (Garnelo et al., 2018). As shown in Table 3, IDFF outperforms the baselines by a large margin, indicating its great potential for simulating molecular dynamics from scratch.
@@ 5.2.2 Sea Surface Temperature Forecasting
This experiment focuses on forecasting sea surface temperature (SST). Based on NOAA OISSTv2, our dataset consists of daily weather images with high-resolution SST data from 1982 to 2021 (Cachay et al., 2023). We created a refined dataset for forecasting SST intervals from 1 to 7 days ahead by strategically subsampling eleven grid tiles of size $60\times 60$ (latitude $\times$ longitude) in the eastern tropical Pacific region (Huang et al., 2021).
We used a class-conditional UNet (ADM) for the SST experiment, with the class encoder part encoding the long-range time dependencies (i.e., for encoding the days) to allow continuous forecasting using the UNet (ADM). See Appendix A.1.4 for the experiment’s network structure and hyperparameter settings. We used the training 1 and sampling 2 algorithms suggested for time-series data. We evaluated the performance of IDFF using the best validation continuous ranked probability score (CRPS) (Matheson and Winkler, 1976) and mean squared error (MSE) on forecasts up to 7 days. CRPS is a proper scoring rule and a popular metric in probabilistic forecasting literature (Gneiting and Katzfuss, 2014). We calculate the CRPS by generating a 20-member ensemble and the MSE on the ensemble mean predictions in Table 4. To assess IDFF’s performance, we compared it against various baseline approaches, including DDPM, DDPM with enabled dropout (Dropout) (Gal and Ghahramani, 2016), DDPM with random perturbation for the initial values (Perturbation) (Pathak et al., 2022), MCVD (Voleti et al., 2022), Dyffusion (Cachay et al., 2023), and OT-CFM (with $NFE=40$) as shown in Table 4. The results for all the baselines are with $NFE=1000$. These baselines provide a robust framework for evaluating the IDFF framework in time-series forecasting. For more details about the SST dataset and experiment, refer to (Cachay et al., 2023). The results show that IDFF (with $NFE=5$) outperforms the baselines in SST forecasting across both metrics, demonstrating great potential for weather forecasting applications. See samples of the forecasting results in Figure 13.
@@
6 Conclusion
IDFF models represent a significant advancement in generative modeling, particularly for time series data, by offering superior computational efficiency compared to traditional flow models like CFMs. The enhanced efficiency and favorable trade-offs between computational cost and sample quality make IDFF a top choice for tasks requiring efficient and high-quality sample generation. Furthermore, IDFF has the potential to extend to other domains, such as music generation (Briot et al., 2017), speech synthesis (Yu and Deng, 2016), and the generation of noisy biological time-series data (Anumanchipalli et al., 2019; Golshan et al., 2020).
One limitation of IDFF pertains to the definition of OT plans, where the computational cost escalates with batch size, resulting in longer training times, especially for large batches. Additionally, while the ADM structure was primarily employed in our experiments, generating large-scale time series data, such as videos, could benefit from utilizing more recent backbone structures as proposed in (Ho et al., 2022; Peebles and Xie, 2023) to enhance efficiency further.
Broader Impacts and Limitations.
IDFF models offer a computationally efficient alternative to traditional flow models like CFMs, enabling faster generation of high-quality samples. This efficiency makes them practical for various applications, suggesting that IDFF could become a preferred choice for generative modeling tasks that demand both efficiency and high sample quality at a lower computational cost. However, it is crucial to establish safeguards and incentives to prevent the misuse of this technology. Care must be taken to ensure that the synthetic data generated by this technology is used following societal norms and ethical standards.
Moreover, IDFF could have broader implications across various fields and should be considered a generic framework that can extend applications such as music generation (Briot et al., 2017), speech generation (Yu and Deng, 2016), and other forms of noisy biological time-series data (Rezaei et al., 2021, 2023; Anumanchipalli et al., 2019; Golshan et al., 2020; Gracco et al., 2005).
@@ 63. Answer: [Yes]
  64. Justification: The experiment section provides details of the datasets and experiments. Additionally, model details are also provided in appendix A.1.4.
@@ A.1.1 Continuity proof for Lemma 1
By defining a vector field as follows:
$\tilde{\bm{u}}_{t}({\mathbf{x}})=\sqrt{(1-\sigma_{t}^{2})}\hat{{\mathbf{v}}}_{t}({\mathbf{x}})+\sigma_{t}^{2}\bm{\xi}_{t},\quad\bm{\xi}_{t}=\gamma\nabla\log p_{t}({\mathbf{x}}_{t})$
where $\gamma$ is a coefficient that controls the contribution of the momentum term $\bm{\xi}_{t}$, we consider a Stochastic Differential Equation (SDE) of the standard form:
$d{\mathbf{x}}_{t}=\tilde{{\mathbf{v}}}_{t}({\mathbf{x}})\,dt+\sigma_{t}\,d{\mathbf{w}}$
with time parameter $t$, drift $\tilde{{\mathbf{v}}}_{t}$, diffusion coefficient $\sigma_{t}$, and $d{\mathbf{w}}$ representing the Wiener process. The solution ${\mathbf{x}}_{t}$ to the SDE is a stochastic process with a probability density, $p_{t}({\mathbf{x}}_{t}|{\mathbf{x}}_{1})$, characterized by the Fokker-Planck equation:
$\frac{dp_{t}({\mathbf{x}}_{t})}{dt}=-\mathrm{div}(\tilde{{\mathbf{v}}}_{t}({\mathbf{x}})p_{t}({\mathbf{x}}_{t}))+\frac{\sigma_{t}^{2}}{2}\Delta p_{t}({\mathbf{x}}_{t})$
where $\Delta$ represents the Laplace operator in ${\mathbf{x}}$. Rewriting this equation in the form of the continuity equation can be done as follows:
$\frac{dp_{t}({\mathbf{x}}_{t})}{dt}$ $=-\mathrm{div}\left(\tilde{{\mathbf{v}}}_{t}({\mathbf{x}})p_{t}({\mathbf{x}}_{t})-\frac{\sigma_{t}^{2}}{2}\frac{\nabla p_{t}({\mathbf{x}}_{t})}{p_{t}({\mathbf{x}}_{t})}p_{t}({\mathbf{x}}_{t})\right)$
$=-\mathrm{div}\left(\left(\tilde{{\mathbf{v}}}_{t}({\mathbf{x}})-\frac{\sigma_{t}^{2}}{2}\nabla\log p_{t}({\mathbf{x}}_{t})\right)p_{t}({\mathbf{x}}_{t})\right)$
$=-\mathrm{div}\left({\mathbf{w}}_{t}p_{t}({\mathbf{x}}_{t})\right)$
where the vector field
${\mathbf{w}}_{t}=\tilde{{\mathbf{v}}}_{t}({\mathbf{x}})-\frac{\sigma_{t}^{2}}{2}\nabla\log p_{t}({\mathbf{x}}_{t})=\sqrt{(1-\sigma_{t}^{2})}\hat{{\mathbf{v}}}_{t}({\mathbf{x}})+\frac{(2\gamma-1)}{2}\sigma_{t}^{2}\nabla\log p_{t}({\mathbf{x}}_{t})$
satisfies the continuity equation with the probability path $p_{t}$ for the equation7.

### v2

@@ 3.3 IDFF time-series adaptation
Input: dataset distribution $p_{1}({\mathbf{x}}_{1})$, initial distribution $p_{0}({\mathbf{x}}_{0})$, bandwidth $\sigma_{0}$, a set of positive weights $\lambda(.)$ and $\beta(.)$, initialized networks $\hat{{\mathbf{x}}}_{1}(.;\theta)$ and $\hat{\bm{\epsilon}}(.;\theta)$,
@@
    ${\mathbf{x}}_{1}\sim p_{1}({\mathbf{x}}_{1})$, ${\mathbf{x}}_{0}\sim p_{0}({\mathbf{x}}_{0})$
@@
    ${\mathbf{x}}_{t}\sim\mathcal{N}(\bm{\mu_{t}},\sigma_{t}^{2}\textbf{I})$
    $\mathcal{L}_{\text{IDFF}}(\theta)\leftarrow\beta(t)^{2}\|\hat{{\mathbf{x}}}_{1}({\mathbf{x}}_{t},t,k;\theta)-{\mathbf{x}}_{1}\|^{2}+\lambda(t)^{2}\left\|\bm{\epsilon}({\mathbf{x}}_{t},t;\theta)-\nabla_{x}\log p_{t}({\mathbf{x}}_{t}|{\mathbf{x}}_{1})\right\|^{2}$
@@
Input: $\hat{{\mathbf{x}}}_{1}(.;\theta)$ and $\hat{\bm{\epsilon}}(.;\theta)$, bandwidth $\sigma_{0}$, time step size $\Delta t$, and $\gamma$.
@@
    ${\mathbf{w}}_{t}\leftarrow\sqrt{(1-\sigma^{2}_{t})}\frac{\hat{{\mathbf{x}}}_{1}({\mathbf{x}}_{t},t;\theta)-{\mathbf{x}}_{t}}{1-t}+\frac{2\gamma-1}{2}\sigma_{t}\hat{\epsilon}({\mathbf{x}}_{t},t;\theta)$
    ${\mathbf{x}}_{t+\Delta t}\sim\mathcal{N}({\mathbf{x}}_{t}+{\mathbf{w}}_{t}\Delta t,\sigma_{t}^{2}\Delta t\textbf{I})$
@@
return Samples $\{{\mathbf{x}}_{1}\}$
@@ 4 Related Work
Recent advancements in CFMs include Schrödinger bridge-based methods (Richter et al., 2023) and conjugate/splitting-based integrators (Pandey et al., 2023), both aimed at improving sampling efficiency. However, these methods still struggle to significantly reduce the NFEs, which remains a critical factor for computational efficiency.
In diffusion models, denoising diffusion implicit models (DDIMs) (Song et al., 2020b) were introduced to address the issue of large NFEs by offering a more efficient sampling process. DDIMs deterministically generate samples from latent variables, unlike DDPMs, which rely on Langevin dynamics. By employing a variational approach, DDIMs accelerate sampling while preserving output quality. Other approaches to enhance sampling in diffusion-based generative models include distillation frameworks such as Flash (Kohler et al., 2024), which generate diverse samples in fewer steps by mitigating training-inference discrepancies.
A distinguishing feature of IDFF is its ability to jointly learn an implicit flow from the generated samples alongside the scoring model. This sets it apart from models like DDIM, which focuses solely on learning a scoring function. By leveraging OT to couple samples between $t=0$ and $t=1$, IDFF accelerates convergence during training while maintaining a low NFE and improving sample quality.
Our image generation experiments, presented in Tables 1 and 2, demonstrate IDFF’s superior performance. Compared to DDPM (Ho et al., 2020), DDIM (Song et al., 2020b), and DPM-Solver (Lu et al., 2022a) (designed to reduce NFEs in diffusion models), IDFF consistently outperforms these models in both efficiency and quality.
@@
In summary, IDFF integrates conditional flow and score-based learning strengths, offering a novel and efficient framework for generative modeling that significantly reduces computational costs while producing high-quality samples. The following section provides more details about these comparisons.
5 Experiments
We conducted two general classes of experiments to demonstrate the effectiveness of IDFF in generative modeling on static data, such as images and time-series data, including simulated chaotic systems (e.g., attractors), molecular dynamics, and sea surface temperature forecasting. The anonymous code for our implementation is available at [this link](https://anonymous.4open.science/r/IDFF-D2E5/README.md).
@@ 5.1 Image Generation
IDFFs exhibit superior performance compared to CFMs in image generation. We trained IDFF models on datasets such as CIFAR-10 and CelebA (additional results for ImageNet-64, CelebA-HQ, LSUN Bedrooms, and LSUN Church are available in Appendix A.5). We utilized a CNN-based UNet (Dhariwal and Nichol, 2021) to simultaneously model both $\hat{{\mathbf{x}}}_{1}(.,t;\theta)$ and $\bm{\epsilon}(.,t;\theta)$ from Equation 11 for all image data except ImageNet-64. For ImageNet-64, we employed a DiT architecture (’DiT-L/4’)(Peebles and Xie, 2022). To implement this, we doubled the input channels of the networks and fed the augmented input, $({\mathbf{x}}_{t},{\mathbf{x}}_{t})$, into the network. The outputs were then split into $(\hat{{\mathbf{x}}}_{1}(.,t;\theta),\bm{\epsilon}(.,t;\theta))$.
@@
To ensure a fair comparison with other models, we used commonly employed metrics to evaluate CFMs’ performance, including negative log-likelihood (NLL) using equation 12 measured in bits per dimension (BPD) Lipman et al. (2022), sample quality as measured by the Frechet Inception Distance (FID), and the average NFE required for the reported FID and NLL, averaged over 50k samples, to assess the computational efficiency of our method. We generate 50k images to calculate the FID against the corresponding reference statistics of each dataset. All models use the same architecture, with specific details provided in Appendix A.4.
We also examined the effect of $\sigma_{0}$ on sample quality for both CIFAR-10 and CelebA datasets in Appendix A.2, finding that a value of $\sigma_{0}=0.6$ achieved the best FID for both datasets, as shown in Table 1 and Table 2. IDFF achieves comparable FID scores with ten times fewer NFEs than CFM and DDPM on the CIFAR-10 and CelebA datasets. IDFF’s qualitative results are provided in Figure 1 (additional samples are available in Appendix A.5).
@@ 5.2.1 Molecular Dynamics Simulation
IDFFs can predict dynamics for complex combinatorial structures such as molecules. Molecular dynamics simulation is a crucial tool in quantum mechanics for understanding the dynamics of molecular behavior at the atomic level (Cazorla and Boronat, 2017). Here, we simulated a fully extended polyalanine structure for 400 picoseconds in a vacuum environment at a temperature of $T=300K$. This molecular structure consists of 253 atoms with 46 dihedral angles. We trained an IDFF model to accurately generate the dihedral angles for the polyalanine structure from scratch. The distributions of the actual and generated dihedral angles for all 46 angles are presented in Figure 3 (A) and (B), respectively. Figure 3 (C) illustrates the dihedral angles for one of the alanine molecules and its associated dihedral angles. A complete trajectory showcasing a pair of actual and generated dihedral angles is depicted in Figure 3 (D).
To assess the performance of IDFF in the MD simulation, we utilized several evaluation metrics, including root mean squared error (RMSE), mean absolute error (MAE), and correlation coefficients (CC) between the generated and actual trajectories of dihedral angles. We benchmarked IDFF against well-known dynamical models for time series, including dynamical VAE models such as SRNNs (Fraccaro et al., 2016), VRNNs (Chung et al., 2015), and Neural ODEs (NODEs) (Garnelo et al., 2018). As shown in Table 3, IDFF outperforms the baselines by a large margin, indicating its potential for simulating molecular dynamics from scratch.
@@ 5.2.2 Sea Surface Temperature Forecasting
IDFFs can predict spatiotemporal dynamics for sea-surface temperature. This experiment centers on forecasting sea surface temperature (SST) (Worsfold et al., 2024). Using the NOAA OISSTv2 dataset, which contains daily high-resolution SST images spanning from 1982 to 2021 (Cachay et al., 2023), we partitioned the data to predict SST intervals from 1 to 7 days ahead. To achieve this, we strategically subsampled eleven grid tiles of size $60\times 60$ (latitude $\times$ longitude) from the eastern tropical Pacific region (Huang et al., 2021).
We used a class-conditional UNet for the SST experiment, with the class encoder part encoding the long-range time dependencies (i.e., for encoding the days) to allow continuous forecasting using the UNet. See Appendix A.4 for the experiment’s network structure and hyperparameter settings. We used the training 4 and sampling 3 algorithms suggested for time-series data. We evaluated the performance of IDFF using the best validation continuous ranked probability score (CRPS) (Matheson and Winkler, 1976) and mean squared error (MSE) on forecasts up to 7 days. CRPS is a proper scoring rule and a popular metric in probabilistic forecasting literature (Gneiting and Katzfuss, 2014). We calculate the CRPS by generating a 20-member ensemble and the MSE on the ensemble mean predictions in Table 4.
To assess IDFF’s performance, we compared it against various baseline approaches, including DDPM, DDPM with enabled dropout (Dropout) (Gal and Ghahramani, 2016), DDPM with random perturbation for the initial values (Perturbation) (Pathak et al., 2022), MCVD (Voleti et al., 2022), Dyffusion (Cachay et al., 2023), and OT-CFM (with $NFE=40$) as shown in Table 4. The results for all the baselines are with $NFE=1000$. These baselines provide a robust framework for evaluating the IDFF framework in time-series forecasting. The results show that IDFF (with $NFE=5$) outperforms the baselines in SST forecasting across both metrics, demonstrating great potential for weather forecasting applications. See samples of the forecasting results in Figure 13.
@@
6 Conclusion and Discussion
IDFF models represent a significant advancement in generative modeling, particularly for time series data. By offering superior computational efficiency compared to traditional flow models like CFMs, IDFF achieves a favorable balance between computational cost and sample quality. This makes IDFF a compelling choice for tasks requiring efficient and high-quality sample generation. There are myriad future possibilities for the application of IDFF in domains such as music generation (Briot et al., 2017), speech synthesis (Yu and Deng, 2016), and modeling biological time-series data (Anumanchipalli et al., 2019; Golshan et al., 2020; Rezaei et al., 2021, 2023; Gracco et al., 2005).
Limitations and Future Work: Computing the OT map for large datasets can be challenging. To address this, we use a minibatch approximation of OT; however, increasing the batch size results in longer training times for larger batches. Second, our experiments primarily utilized a CNN-based UNet structure (Dhariwal and Nichol, 2021) as the backbone of the models. We anticipate that some of these limitations may be mitigated by utilizing new neural network structures (Peebles and Xie, 2023).
Broader Impacts: The computational efficiency of IDFF models makes it attractive for use in applications where speed and high sample quality are crucial, e.g., in resource-constrained environments. As with any powerful generative technology, the ethical implications must be considered. Safeguards and incentives should be established to prevent misuse of this technology. The synthetic data generated using IDFF must be used following societal norms and ethical standards.
@@ 63. Answer: [Yes]
  64. Justification: The experiment section provides details of the datasets and experiments. Additionally, model details are also provided in appendix A.4.
@@ A.1.1 Continuity proof for Lemma 1
We define a vector field as follows:
$\tilde{\mathbf{v}}_{t}(\mathbf{x}_{t})=\sqrt{1-\sigma_{t}^{2}}{\mathbf{v}}_{t}(\mathbf{x}_{t})+\sigma_{t}^{2}\bm{\xi}_{t},\quad\bm{\xi}_{t}=\gamma\nabla\log p_{t}(\mathbf{x}_{t})$ (13)
where $\gamma$ is a coefficient controlling the contribution of the momentum term $\bm{\xi}_{t}$.
Consider a Stochastic Differential Equation (SDE) of the standard form:
$d\mathbf{x}_{t}=\tilde{\mathbf{v}}_{t}(\mathbf{x}_{t})\,dt+\sigma_{t}\,d\mathbf{w}$ (14)
with time parameter $t$, drift $\tilde{\mathbf{v}}_{t}$, diffusion coefficient $\sigma_{t}$, and $d\mathbf{w}$ representing the Wiener process.
The solution $\mathbf{x}_{t}$ to the SDE is a stochastic process with probability density $p_{t}(\mathbf{x}_{t}|\mathbf{x}_{1})$, characterized by the Fokker-Planck equation:
$\frac{\partial p_{t}(\mathbf{x}_{t})}{\partial t}=-\nabla\cdot(\tilde{\mathbf{v}}_{t}(\mathbf{x}_{t})p_{t}(\mathbf{x}_{t}))+\frac{\sigma_{t}^{2}}{2}\Delta p_{t}(\mathbf{x}_{t})$ (15)
where $\Delta$ represents the Laplace operator in $\mathbf{x}_{t}$.
We can rewrite this equation in the form of the continuity equation:
$\frac{\partial p_{t}(\mathbf{x}_{t})}{\partial t}$ $=-\nabla\cdot\left(\tilde{\mathbf{v}}_{t}(\mathbf{x}_{t})p_{t}(\mathbf{x}_{t})-\frac{\sigma_{t}^{2}}{2}\frac{\nabla p_{t}(\mathbf{x}_{t})}{p_{t}(\mathbf{x}_{t})}p_{t}(\mathbf{x}_{t})\right)$ (16)
$=-\nabla\cdot\left(\left(\tilde{\mathbf{v}}_{t}(\mathbf{x}_{t})-\frac{\sigma_{t}^{2}}{2}\nabla\log p_{t}(\mathbf{x}_{t})\right)p_{t}(\mathbf{x}_{t})\right)$ (17)
$=-\nabla\cdot\left(\mathbf{w}_{t}p_{t}(\mathbf{x}_{t})\right)$ (18)
where the vector field $\mathbf{w}_{t}$ is defined as:
$\mathbf{w}_{t}(\mathbf{x}_{t})=\tilde{\mathbf{v}}_{t}(\mathbf{x}_{t})-\frac{\sigma_{t}^{2}}{2}\nabla\log p_{t}(\mathbf{x}_{t})=\sqrt{1-\sigma_{t}^{2}}{\mathbf{v}}_{t}(\mathbf{x}_{t})+\frac{2\gamma-1}{2}\sigma_{t}^{2}\nabla\log p_{t}(\mathbf{x}_{t})$ (19)
As $t$ approaches 0 or 1, $\sigma_{t}\rightarrow 0$, causing $\tilde{\mathbf{v}}_{t}(\mathbf{x}_{t})$ to converge to $\hat{\mathbf{v}}_{t}(\mathbf{x}_{t})$. This ensures that the continuity equation is satisfied with the probability path $p_{t}(\mathbf{x}_{t})$ for the original equation.